OpenStudy (anonymous):

10=the square root of (2x +2) +5

OpenStudy (saifoo.khan):

\[10 = \sqrt{2x+2} + 5\]

OpenStudy (saifoo.khan):

\[\sqrt{2x+2}= 5\]

OpenStudy (saifoo.khan):

2x+2 = 25 Solve for x.

OpenStudy (anonymous):

ok thanks :)